### v1

# EgoTextVQA: Towards Egocentric Scene-Text Aware Video Question Answering

Abstract
[…]
Figure 1: Our EgoTextVQA aims for QA assistance involving scene text from an ego-perspective mainly in outdoor driving (EgoTextVQA-Outdoor) and indoor house-keeping (EgoTextVQA-Indoor), with the questions reflecting the real user needs yet without the visual focus on scene text. Benchmarking results show that all models struggle on EgoTextVQA, highlighting continued efforts for improvements.
footnotetext: Work done when the first author was a visiting student at NUS.footnotetext: $\dagger$ Corresponding author.
1 Introduction
The advances in wearable cameras and egocentric vision research [15, 26, 37], has led to a surge of interest in developing assistance applications. A promising direction is egocentric video question answering (VideoQA) [13, 52, 6, 12, 56], where the AI agents offer live help by answering user questions. Despite the potential, related dataset researches [54, 13, 20, 31, 10] have been designed to assess the VideoQA models in their visual understanding capabilities. The questions are not curated to reflect real-user needs in assistive applications (see EgoTaskQA in Fig. 1). QAEgo4D [2] and AssistQ [50] are closer in aim, but these datasets are designed to assist with episodic memory and “how to” demonstration; they emphasize pure visual understanding while neglecting scene text.
Research in assistive techniques [16, 42] have shown that scene text, including signs, labels, and other text elements that are ubiquitous in our daily lives, is frequently involved in user questions when seeking assistance. Existing efforts for QA on scene-text [42, 3] primarily aim at Optical Character Recognition (OCR) from ideal imaging conditions. They assume that people take good pictures and ask questions directly pinpointing the scene-text regions (see TextVQA [42] in Fig. 1). Despite the progress (e.g., leading results have accuracies of 85% on ST-VQA [3, 47]), such a simplified setting has limited application value. Consider, for example, those with a visual impairment [16], they will likely struggle to take clean and in-focus images; or can they point to scene-text regions. More recent works [61, 45] begin to study text-based VideoQA which allows QA on multiple images to reduce the proportion of unanswerable cases. Nonetheless, they still assume people know well the locations of scene text (see RoadTextVQA [45] in Fig. 1). Corresponding VQA models, while achieving good results on such benchmarks, cannot reason over user intentions other than an extraction of the OCR results [19].
In light of this, we introduce EgoTextVQA, a novel and rigorously constructed Egocentric Scene-Text Aware Video Question Answering benchmark. EgoTextVQA is designed to advance research on egocentric QA assistance in real-life scenarios. We collect 1.5K ego-view videos and 7K scene-text aware questions from outdoor driving (EgoTextVQA-Outdoor) and indoor house-keeping (EgoTextVQA-Indoor) activities. By emphasizing real user needs, we allow questions and answers that do not explicitly pinpoint the scene text. However, comprehending scene text is crucial to answering the questions (see examples in Figure 1). Moreover, to simulate real-time and streaming video QA, we set a timestamp for each question and allow the models to access only visual contents before the question’s timestamp to make related responses. Detailed visual scenarios and question types are also provided for better analysis of models.
For its realistic setting, EgoTextVQA retains all the challenges in existing scene text VQA datasets and incorporates new major difficulties: First and foremost, it is crucial to reason about the user’s intentions with respect to the visual scene and scene text to both understand and to answer the questions. Second, the model may need to reason across multiple frames to either infer the user’s behavior or to locate the required scene text for answering. Third, the model may need to infer the user’s current state when posing the questions and provide actionable answers for meaningful assistance. Finally, the egocentric dynamic situation poses an additional challenge to scene text recognition as opposed to the well-captured images and focused scene text.
With EgoTextVQA, we comprehensively benchmark 10 prominent models that perform well on existing scene-text VQA datasets, and find they all struggle, especially the open-source ones like MiniCPM-V 2.6 [53], ShareGPT4Video [7] and CogVLM2-Video [18]. The best-performing model is Gemini 1.5 Pro [39], yet its accuracy is still lower at around 33%. Comprehensive analyses and additional heuristic explorations reveal many insights on model behaviors and possible directions for improvements:
  1. 1.
    Temporally localizing the key frames and jointly reasoning over multiple key frames (vs. a single key frame) are crucial for improvements, especially for long video QA scenarios on EgoTextVQA-Indoor.
  2. 2.
    High-resolution image and scene text input can significantly boost models’ performances. Yet, one needs to mind the compute efficiency. Also, models that inherently take fixed higher-resolution inputs are not necessarily better than those using lower-resolution inputs.
  3. 3.
    Scene text from additional OCR techniques, despite being auxiliary, are extremely helpful for all models.
Additionally, we find that even humans cannot perform well on this task (43% on EgoTextVQA-Outdoor and 27% on EgoTextVQA-Indoor), largely due to the difficulty of scene text recognition and the limited knowledge of humans, highlighting the significance of this research. To summarize our contributions: 1) We propose to study scene-text aware VideoQA towards egocentric assistance and construct the EgoTextVQA dataset containing both EgoTextVQA-Indoor and EgoTextVQA-Outdoor. 2) We benchmark 10 contemporary powerful models covering both open-source and closed-source ones and comprehensively analyze their limitations. 3) We explore a series of heuristic methods for improvements and share insightful findings for future work.
2 Related Work
2.1 Scene-Text VQA Benchmark
In the scene-text VQA field, various image- and video-based datasets [42, 49, 3, 61, 45, 34, 28] have been proposed. The datasets like TextVQA [42], ST-VQA [3], and ESTVQA [49] provide high-quality images and questions that clearly point to scene text, but such simple settings limit practical applications. To expand the application scope, the datasets like M4viteVQA [61] and RoadTextVQA [45] offer text-rich videos, but design simple questions referring to well-focused scene text. Such simplified settings make QA less practical and they challenge no more than an identification of the scene text. In addition, RoadTextVQA focuses merely on road driving. In this work, we introduce EgoTextVQA to advance research in real-life QA assistance from an ego-perspective. Our videos cover diverse daily scenarios, with the questions reflecting real-user needs yet without the visual attention on scene text.
2.2 MLLMs for Scene-Text VQA
Recent advancements in Multimodal Large Language Models (MLLMs) [60, 48, 57] have shown significant potential in addressing scene-text VQA tasks [42, 3, 28]. Current MLLMs [29, 60, 23] achieve substantial performance gains in image scene-text QA by enhancing their ability to process higher-resolution images, leading to significant improvements and strong performance. However, these models are limited to image-level input, leaving their video comprehension capabilities largely unexplored. To address this limitation, recent models [53, 58, 47, 25, 8, 18, 7] are developed to support video-level or multi-frame inputs while retain their performance on image-based scene-text QA datasets. Nonetheless, these models have been evaluated separately on image-based OCR-rich benchmarks [42, 3, 28, 34, 32, 33] and video understanding benchmarks [31, 51, 14, 24], lacking suitable assessment for scene-text comprehension in more realistic and dynamic environments. In this work, we will comprehensively analyze the model behaviors in egocentric QA assistance in real-life dynamic environments.
3 EgoTextVQA Dataset
Figure 2: Examples of EgoTextVQA. Scene text plays pivotal role in understanding and answering the questions which reflect real-user needs. Yet, the videos are without the visual focus on scene text.
3.1 Dataset Creation
We leverage existing text-rich video sequences and the multimodal understanding and generation capabilities of MLLMs to create our dataset. The generated QA candidates are carefully checked and corrected by human annotators as necessary to ensure the dataset’s high diversity and quality.
Table 1: Data distribution and related examples of EgoTextVQA. The green words denote scene texts that appear in the video.
  Category #Q / #V QA Examples
            EgoTextVQA-Outdoor
Location 1,505/553 Q: Where should I go if I need a dentist? A: Vijaya Dental Hospital.
Description 1,387/601 Q: What is the license plate of the car in front of me? A: Ts07ge5554.
Direction 952/468 Q: Which direction do I go to get to Vapor Inn? A: Continue straight and turn right at the next road.
Intention Reasoning 813/443 Q: What should I be cautious of in this area? A: Watch for pedestrians crossing at the crosswalk.
Others 191/152 Q: What is the main theme of the event advertised on the bus and when and where is it happening? A: The main theme is “Race Against HIV”, and it is happening on Dec 1, 2018, at People’s Plaza.
EgoTextVQA-Indoor
Hands-on 698/284 Q: How should I properly handle the AVT equipment? A: Maintain a firm grip and use proper posture.
Shopping 258/78 Q: Where can I find the “Milano” cookies? A: Top shelf on the right.
Kitchen 335/149 Q: What should I use to clean up countertop spills? A: Ecover dish soap
Book-related 323/103 Q: Where can I find information on learning the guitar? A: Guitar for beginners.
Gameplay 301/124 Q: What card did my opponent play before I placed the red 9? A: Red 8.
Others 201/88 Q: Where might I find the office rules? A: On the notices on the wall near the entrance.
Raw Video Filtering Our videos are drawn from two public ego-view video datasets, RoadTextVQA [45] and EgoSchema [31]. They cover outdoor and indoor scenarios, centered around driving and housekeeping activities. RoadTextVQA has 3,222 ten-second videos from dashcam footage of people driving. EgoSchema features 5,063 three-minute videos sourced from Ego4D [15]; its videos are mostly about indoor activities such as cooking, game playing, manufacturing, etc. To find videos rich in scene text, we first apply a state-of-the-art scene-text detection system [17] to the raw videos and threshold for videos with a significant percentage of frames with scene text: exceeding 5% for Ego4D and 15% for RoadTextVQA (since it features more scene text). After automatic filtering, we obtain 700 videos from RoadTextVQA [45] and 1,800 from EgoSchema [15]. Since Ego4D videos are not collected for scene text study, we further filter the unqualified videos, mainly removing the tedious videos and videos that contain watermark or text present only on the clothes of the camera wearers. Finally, we obtain 933/700 videos from Ego4D/RoadTextVQA.
QA Generation Manually collecting QA pairs is labor-intensive; the QAs may also lack diversity. We exploit advanced MLLMs (i.e., GPT-4o) and generate the initial QAs with the following protocol: First, a video is decoded into 6 frame-per-second (fps) and evenly divided into 5 segments, wherein the frames without scene text are removed. Then, we uniformly sample 3 frames from each segment, and feed them to GPT-4o along with the prompts (in Appendix D) to generate 3 QA pairs. This leads to 15 or fewer QA pairs for each video. The prompts are carefully designed to elicit questions that are: (1) goal-oriented, capturing real user needs by engaging with the visual scene from a first-person perspective; (2) aware of scene text, though not necessarily requiring the exact text transcription; (3) naturally expressed in a colloquial, first-person manner with referring expressions; and (4) challenging, as they demand an understanding of the video beyond a single image.
Figure 3: Distribution of QAs and OCR numbers.
Table 2: Comparison of related egocentric VQA and scene-text VQA test benchmarks. ST: Scene Text. STQ/STA: Questions/Answers contain scene text. EgoV: Egocentric Video. IntQ: Intentional Questions. QC: Question Category. OE/MC: Open-ended/Multi-choice.
  Benchmark #Que. #Video Ave. Len. (s) ST STQ (%) STA (%) EgoV Indoor Outdoor Real Time IntQ QC Task
      Egocentric VQA Benchmarks
EgoThink [11] 700 - - - - OE
QAEgo4D [2] 1,854 166 495.1 - - OE
AssistQ [50] 531 100 115.0 - - MC
EgoSchema [31] 5,063 5,063 180.0 - - MC
EgoMemoria [54] 7,026 629 858.5 - - MC
Scene-Text VQA Benchmarks
TextVQA [42] 5,734 - - 16 85 OE
ST-VQA [3] 4,163 - - 18 92 OE
ESTVQA [49] 5,014 - - 25 97 OE
M4ViteVQA [61] 2,103 680 5.7 30 96 OE
RoadTextVQA [45] 1,052 329 10.0 60 65 OE
EgoTextVQA (Ours) 7,064 1,507 101.7 52 45 OE
Manual Participation Despite using well-designed prompts, the generated QA pairs may fail to meet our strict requirements. Some generated questions are hallucinated or irrelevant to the visual content, have wrong formats, or are gibberish, e.g., involving special typeset characters not part of spoken language. We first employ GPT-4o to answer the questions without videos and remove the questions that can be answered blindly ($\sim$10%). We then conduct manual filtering and correction for the remaining QAs. We provide some representative generated QA errors in Appendix A.
Specifically, we invite 9 well-trained university students for annotation. The manual check has 5 stages: (1) five annotators review for problematic QAs which may be redundant, not reflect real-user needs, and irrelevant to scene text. Where possible, the questions are amended; otherwise, they are removed, leaving half of the original QAs; (2) and (3) two additional annotators re-examine and correct any mechanical issues, further reducing the QAs to 30%; (4) five annotators refine the data to improve its colloquial quality; (5) four annotators strictly refine the remaining QAs according to the criteria specified in QA generation. The participants at each stage are asked to first check the remaining QAs from the prior stage. Additionally, we manually enrich $\sim$10time-sensitive questions to simulate real-time live QA by copying each suitable question to a different timestamp and invite five additional annotators to provide concise answers based on the visual content at the new timestamp. Ultimately, we obtain 7,064 QA pairs of 1,507 videos to form the EgoTextVQA dataset. Detailed question type and video scenario labels are additionally provided by the authors. Figure 2 shows some representative examples; more are presented in Appendix B.5.
3.2 Dataset Analysis
Dataset Statistics EgoTextVQA characterizes assistance-seeking questions related to scene text in diverse real-life scenarios. It consists of two parts: (1) EgoTextVQA-Outdoor focuses on the outdoor scenarios, with 694 videos and 4,848 QA pairs that may arise when driving; (2) EgoTextVQA-Indoor  emphasizes indoor scenarios, with 813 videos and 2,216 QA pairs that users may encounter in house-keeping activities. Distributions of videos and questions are listed in Table 1. We also analyze the most frequent question and answer patterns in Figure Fig. 3 ①, ②. The statistics reveal a clear discrepancy between QAs; the driving questions are mostly related to navigation which pure scene text cannot answer, while the house-keeping questions are information seeking, where answers can be extracted from scene text. Figure 3 ③ shows that the EgoTextVQA-Indoor videos have a long-tailed distribution in the number of OCR tokens, with the majority having less than 200. In contrast, EgoTextVQA-Outdoor videos are more evenly distributed with around 500.
Dataset Comparison We compare EgoTextVQA with several related benchmarks in Table 2. EgoTextVQA has several unique features. Firstly, it is the first VideoQA testbed designed for egocentric scene-text aware QA assistance in the wild, containing 7K QA pairs that reflect diverse user intentions across 1.5K different egocentric visual scenarios covers both indoor and outdoor activities. The QA pairs focus on scene text comprehension, with approximately 45% of answers and 52% of questions directly referencing the exact scene text. Notably, EgoTextVQA is also the first VideoQA dataset to introduce real-time QA, providing detailed timestamps for each question. While EgoMemoria [54] also includes question timestamps, its answers are based on the entire video content. In contrast, the answers in EgoTextVQA are derived from the video content before the question timestamp and vary according to the specific timestamp. Moreover, the questions in EgoTextVQA emphasize user intention inferring rather than pure visual understanding, and question categories are provided to facilitate model analysis.
4 Experiments
4.1 Experimental Setups
Table 3: Performance of SoTA MLLMs on EgoTextVQA-Outdoor. Res. denotes image resolution. We uniformly sample #F frames that the models can accept from only video content before the timestamp. In this table, we test the performance of the model on low-resolution videos (640$\times$360 and 960$\times$540) of the EgoTextVQA-Outdoor  dataset. The human study is conducted on a random 30% of all questions, with the only constraint that answers rely on video content before the question timestamp. The best and second-best results are highlighted.
  Method LLM Res. #F Location Direction Description Int. Reasoning Others Average
  Human - - - 38.7 / 2.3 32.4 / 2.2 54.9 / 3.0 45.3 / 2.6 36.9 / 2.5 43.1 / 2.6
Open-source Models
ShareGPT4Video [7] LLaMA3-8B - 60 11.7 / 1.5 20.8 / 1.8 5.5 / 0.7 10.7 / 1.3 10.0 / 1.2 11.5 / 1.3
CogVLM2-Video [18] LLaMA3-8B - 60 10.6 / 1.5 15.7 / 1.7 12.4 / 1.0 12.4 / 1.3 10.0 / 1.3 12.4 / 1.4
MiniCPM-V 2.6 [53] Qwen2-7B $448^{2}$ 60 7.7 / 0.5 14.4 / 0.9 15.1 / 0.9 8.7 / 0.6 10.0 / 0.8 11.4 / 0.7
VILA1.5 [25] LLaMA3-8B $384^{2}$ 32 14.6 / 1.5 20.2 / 1.6 18.2 / 1.3 18.3 / 1.5 14.7 / 1.5 17.4 / 1.5
LLaVA-NeXT-Video [58] Qwen2-7B $384^{2}$ 60 18.1 / 1.3 22.5 / 1.5 19.6 / 1.3 19.3 / 1.4 15.7 / 1.4 19.5 / 1.4
InternVL2 [8] InternLM2.5-7B $448^{2}$ 32 15.8 / 1.4 21.9 / 1.7 14.8 / 1.0 14.5 / 1.2 13.6 / 1.3 16.4 / 1.3
InternVL2 [8] InternLM2-20B $448^{2}$ 16 21.5 / 1.7 27.5 / 2.0 23.5 / 1.5 22.9 / 1.7 22.5 / 1.7 23.5 / 1.7
Qwen2-VL [47] Qwen2-7B - 16 23.4 / 1.8 35.2 / 2.4 30.8 / 1.9 25.6 / 1.9 22.5 / 1.8 28.2 / 2.0
Closed-source Models
GPT-4o [35] - $768^{2}$ 32 28.3 / 1.6 24.9 / 1.4 35.0 / 1.9 31.6 / 1.8 29.8 / 1.8 30.3 / 1.7
Gemini 1.5 Flash [39] - $768^{2}$ 32 25.4 / 1.6 28.3 / 1.8 38.1 / 2.2 28.1 / 1.8 26.6 / 2.0 30.1 / 1.9
Gemini 1.5 Pro [39] - $768^{2}$ 32 33.2 / 2.1 28.3 / 1.8 38.8 / 2.2 32.7 / 2.0 34.6 / 2.2 33.4 / 2.0
Evaluation Metric We use GPT-4o mini [36] to evaluate the semantic similarity between the generated and the ground-truth (GT) answers, thus aligning closely with human scoring. Following prior work [30], we assess the generated answers with two metrics: Accuracy (0-100%, the percentages of “yes” answers from evaluator) and Score (0-5, with 5 being the highest match). Concretely, we format the question, predicted answer, and GT answer along with our customized evaluation prompts (in Appendix D) for GPT-4o mini to determine the “yes/no” and score.
Model Evaluation We collect three closed-source API-based models (GPT-4o [35], Gemini 1.5 Flash [39], Gemini 1.5 Pro [39]) and seven advanced open-source MLLMs (MiniCPM-V 2.6 [53], ShareGPT4Video [7], InternVL2 [8], VILA1.5 [25], LLaVA-NeXT-Video [58], CogVLM2-Video [18], Qwen2-VL [47]) for evaluation. The models have reported the state-of-the-art (SOTA) on existing scene-text VQA benchmarks [3, 42, 16], e.g., Qwen2-VL achieves QA accuracy $\sim$85on ST-VQA [3]. More details about these MLLMs and their specific prompts are presented in Appendix B.1 and D.
4.2 Result Analysis
Tables 3 and  4 show that all models, especially the open-source ones, struggle to perform well on EgoTextVQA. The best results (Gemini Pro 1.5 [39]) are around 33-34% on both EgoTextVQA-Outdoor and EgoTextVQA-Indoor. This exceeds the best-performing open-source models by $\sim$5and $\sim$9on EgoTextVQA-Outdoor and EgoTextVQA-Indoor respectively. The performances vary significantly among the open-source models, with Qwen2-VL [47] and LLaVA-NeXT-Video [58] being the best while MiniCPM-V 2.6 [53] and ShareGPT4Video [7] the worst.
Table 4: Performance of SoTA MLLMs on EgoTextVQA-Indoor. Res. denotes image resolution. We uniformly sample the #F frames that the model can accept from all video frames before the timestamp as input. In this table, we test the performance of the model on the videos (the resolution mainly includes 480$\times$360 and 640$\times$360) of the EgoTextVQA-Indoor  dataset. The human study is conducted on a randomly sampled 30% of all questions. The best and second-best Accuracy / Score results are bolded and underlined respectively.
  Method LLM Res. #F Hands-on Kitchen Shopping Gameplay Book-Related Others Average
  Human - - - 26.6 / 1.9 31.3 / 2.1 18.2 / 1.4 34.8 / 2.2 27.2 / 1.8 32.8 / 2.1 27.7 / 1.9
Open-source Models
ShareGPT4Video [7] LLaMA3-8B - 128 7.0 / 1.0 5.1 / 1.0 5.9 / 1.3 12.3 / 1.2 7.1 / 1.1 4.0 / 0.9 7.0 / 1.1
CogVLM2-Video [18] LLaMA3-8B - 128 12.3 / 1.4 11.6 / 1.4 10.9 / 1.4 23.6 / 1.8 11.2 / 1.2 13.4 / 1.4 13.5 / 1.4
MiniCPM-V 2.6 [53] Qwen2-7B $448^{2}$ 120 14.2 / 0.8 8.1 / 0.5 10.3 / 0.6 17.6 / 1.1 16.4 / 1.0 12.9 / 0.8 13.3 / 0.8
VILA1.5 [25] LLaMA3-8B $384^{2}$ 32 16.5 / 1.5 18.8 / 1.5 15.1 / 1.5 24.9 / 1.9 18.9 / 1.6 17.4 / 1.5 18.2 / 1.6
InternVL2 [8] InternLM2.5-7B $448^{2}$ 32 13.5 / 1.2 12.5 / 1.1 14.5 / 1.3 13.3 / 1.2 14.9 / 1.3 12.4 / 1.1 13.6 / 1.2
InternVL2 [8] InternLM2-20B $448^{2}$ 16 20.3 / 1.6 22.4 / 1.8 18.2 / 1.6 25.6 / 1.8 19.2 / 1.6 21.9 / 1.7 21.0 / 1.7
Qwen2-VL [47] Qwen2-7B $448^{2}$ 48 21.5 / 1.8 21.5 / 1.7 22.5 / 1.8 29.9 / 2.1 22.0 / 1.8 26.4 / 1.8 23.3 / 1.8
LLaVA-NeXT-Video [58] Qwen2-7B $384^{2}$ 128 24.4 / 1.7 23.6 / 1.7 21.8 / 1.7 29.9 / 1.9 27.9 / 1.9 27.8 / 1.8 25.4 / 1.8
Closed-source Models
GPT-4o [35] - $768^{2}$ 60 25.4 / 1.6 30.2 / 1.9 29.6 / 1.9 32.9 / 1.9 29.7 / 1.9 23.9 / 1.5 28.3 / 1.8
Gemini 1.5 Flash [39] - $768^{2}$ 60 29.7 / 2.0 33.7 / 2.1 32.3 / 2.1 34.5 / 2.3 33.6 / 2.2 30.1 / 2.0 32.0 / 2.1
Gemini 1.5 Pro [39] - $768^{2}$ 60 33.2 / 2.1 35.2 / 2.1 33.8 / 2.1 39.5 / 2.4 34.3 / 2.1 30.9 / 2.0 34.4 / 2.1
Figure 4: Result Visualization.
EgoTextVQA-Outdoor Table 3 shows that the MLLMs consistently perform poorly across different question categories. The open-source models particularly fall short in answering “Location” and “Intention Reasoning” types of questions, while the closed-source models GPT-4o and Gemini tend to fail on “Direction” questions. “Location” questions require a nuanced understanding of the spatial positions of targets within the visual scenes, while “Intention Reasoning” questions demand strong inference capabilities of user intentions conditioned on the egocentric visual situation. The poor performance on these questions reveals the severe limitations of existing open-source models in fine-grained visual understanding and reasoning in egocentric scenarios. For the poor behavior of closed-source models on “Direction” questions, a careful observation of the failure predictions shows that models often answer like “unknown scene text” due to their inability to locate or recognize specific scene text (see Figure 4 Top). Overall, Gemini 1.5 Pro [39] wins across most categories, reaching an overall accuracy of 33.4% and a score of 2.0, establishing it as the SOTA model on EgoTextVQA-Outdoor. However, its performance is still 10% behind humans. This results show that even for powerful MLLMs, perceiving scene texts in complex, dynamic, ego-perspective environments is still a challenging and open research topic.
In Figure 5, we additionally analyze the model behaviors on the real-time QA subset of EgoTextVQA-Outdoor. The results show that all models perform much worse on this subset, with the highest accuracy being 20.2% (vs. 33.4% on the full set), highlighting the significant challenge of egocentric live QA. Interestingly, MiniCPM-V 2.6 [53] stands among the Top-2 open-source models for real-time QA despite it being the worst on the full set. Nevertheless, a careful observation of predictions shows that MiniCPM primarily excels at answering unanswerable questions with “I did not find it”, compared with other models which tend to hallucinate the answers. This reflects the models are essentially weak in providing meaningful answers to satisfy user needs.
Figure 5: Performance of MLLMs on the real-time QA subset of EgoTextVQA-Outdoor ($\sim$623 QA pairs).
Table 5: Effects of taking as input an image (“TS Frame”) and three images (“QA Frames” analogous to QA generation) at the question timestamp.
  Method Input EgoTextVQA-Outdoor EgoTextVQA-Indoor
Accuracy Score Accuracy Score
  InternVL2-8B [8] Video 16.4 1.3 13.6 1.2
TS Frame 15.8 $\downarrow$ 0.6 1.2 $\downarrow$ 0.1 6.5 $\downarrow$ 7.1 0.7 $\downarrow$ 0.5
QA Frames 18.5 $\uparrow$ 2.1 1.4 $\uparrow$ 0.1 17.7 $\uparrow$ 4.1 1.4 $\uparrow$ 0.2
MiniCPM-V 2.6 [53] Video 11.4 0.7 13.3 0.8
TS Frame 7.9 $\downarrow$ 3.5 0.2 $\downarrow$ 0.5 3.8 $\downarrow$ 9.5 0.3 $\downarrow$ 0.4
QA Frames 9.3 $\downarrow$ 2.1 0.6 $\downarrow$ 0.1 13.7 $\uparrow$ 0.4 0.9 $\uparrow$ 0.1
Qwen2-VL [47] Video 28.2 2.0 23.3 1.8
TS Frame 30.9 $\uparrow$ 2.7 2.1 $\uparrow$ 0.1 15.7 $\downarrow$ 7.6 1.4 $\downarrow$ 0.4
QA Frames 31.9 $\uparrow$ 3.7 2.2 $\uparrow$ 0.2 30.7 $\uparrow$ 7.4 2.2 $\uparrow$ 0.4
GPT-4o [35] Video 30.3 1.7 28.3 1.8
TS Frame 27.7 $\downarrow$ 2.6 1.6 $\downarrow$ 0.1 14.8 $\downarrow$ 13.5 0.9 $\downarrow$ 0.9
QA Frames 30.4 $\uparrow$ 0.1 1.8 $\uparrow$ 0.1 41.6 $\uparrow$ 13.3 2.4 $\uparrow$ 0.6
Gemini 1.5 Pro [39] Video 33.4 2.0 34.4 2.1
TS Frame 30.4 $\downarrow$ 3.0 1.8 $\downarrow$ 0.2 15.8 $\downarrow$ 18.6 1.1 $\downarrow$ 1.0
QA Frames 26.3 $\downarrow$ 7.1 1.6 $\downarrow$ 0.4 38.2 $\uparrow$ 3.8 2.4 $\uparrow$ 0.3
EgoTextVQA-Indoor Table 4 shows that MLLMs have difficulty in answering user questions in various indoor scenarios. Concretely, the open-source models are especially weak in answering questions in “Kitchen” and “Shopping” activities. Our qualitative analysis (see Figure 4 Bottom) shows that people often ask about the location of ingredients and utensils in the kitchen, where the models tend to provide vague positional information (e.g., “on the left” rather than “on the top shelf on the left”). In shopping, it is particularly challenging for the models to locate a user-specified item among large and densely packed products. Interestingly, we find that ShareGPT4Video [7] performs the worst across all scenarios even though it includes Ego4D videos for training, indicating that it may not be a scene-text aware MLLM. Again, the closed-source models perform relatively better across different visual scenarios.
Intriguingly, we find that humans perform even worse than the closed-source models. Feedback from our human annotators and a careful analysis of the answers suggest that the primary reasons are as follows: (1) scene text recognition is challenging for humans in Ego4D videos due to lower resolution, motion blur, occlusion, and long video; (2) answering the questions needs external knowledge beyond simple scene text recognition. EgoTextVQA-Outdoor mainly involves knowledge about store or restaurant chains (e.g., “Home Deposells building materials), and EgoTextVQA-Indoor focuses on items like kitchen utensils, board games, and specialized books (e.g., identifying amandolineslices potatoes or aquantumfield theory book); (3) answer diversity also results in relatively lower human performance. The first two points underscore the challenges and significance of research on egocentric scene-text QA assistance. The third point unveils a limitation of our dataset, which we wish to solve in the feature by enriching the ground-truth answers for each question. We provide more human study analysis in Appendix B.4 and C.
Outdoor Driving vs. Indoor House-Keeping While humans perform worse on EgoTextVQA-Indoor, the MLLMs are basically at the same performance level on both datasets. In particular, we find Qwen2-VL [47] stands out as the best open-source model on EgoTextVQA-Outdoor, whereas LLaVA-NeXT-Video [58] ranks the top on EgoTextVQA-Indoor. We speculate that LLaVA-NeXT-Video is better at coping with long videos. More model design and case analysis are provided in Appendix B.2 and B.5.
5 Heuristic Solution Investigations
To facilitate future study, we conduct extensive experiments to thoroughly analyze the challenges presented by EgoTextVQA in Tables 5 to 7 and Figures 5 to  7. Our analysis centers on the following four questions:
Q1: Can image-level understanding solve our task? Sometimes. We test the models by inputting a single frame at the question timestamp (“TS Frame”). Compared with using uniformly sampled video frames, Table 5 shows that almost all the models have a significant performance drop ranging from 0.6$\sim$3.5on EgoTextVQA-Outdoor and as much as 7.1$\sim$18.6on EgoTextVQA-Indoor. This signifies the importance of video-level reasoning, especially for EgoTextVQA-Indoor, where questions feature longer video action understanding. Notably, Qwen2-VL improves its performance when fed with a single frame on EgoTextVQA-Outdoor. We speculate that Qwen2-VL excels at scene text recognition at static frames but shows insufficient capability in handling information redundancy across multiple frames.
Figure 6: Effects with different numbers of frame inputs sampled at 1fps backward from the question timestamp.
Q2: Is temporal grounding important? Sometimes. We take the three frames used in question generation as model inputs (“QA Frames”). Table 5 shows that this effectively boosts all model performances on EgoTextVQA-Indoor. However, the improvements on EgoTextVQA-Outdoor are not stable among models. Specifically, Qwen2-VL [47] and InternVL2-8B [8] increase their accuracy by 3.7% and 2.1% respectively but MiniCPM-V 2.6 [53] and Gemini 1.5 Pro decrease by 2.1% and 7.1%, respectively. Overall, the improvement suggests there is space for models to optimize their temporal grounding, i.e., to localize the key frames for answering. In contrast, the decrease in performance indicates that the models may be vulnerable to seriously reduced information input. It may also reflect the great efforts made by annotators for correcting mistakes from the automatic QA generation, as also exemplified by the poor results of GPT-4o (QA generator).
We investigate a question-timestamp aware sampling strategy: starting from the question timestamp and sampling backwards at 1 fps. This strategy is inspired by intuition that people often pose questions based on their most recent visual settings. Figure 6 shows that incorporating such a sampling strategy significantly improves Qwen2-VL’s performance on EgoTextVQA-Outdoor, even with fewer frames. However, for Gemini 1.5 Pro, the opposite is true and more frames are key for better performance. On EgoTextVQA-Indoor, we find that both Qwen2-VL and Gemini 1.5 Pro shrink their performances compared with standard uniform sampling across the whole video, possibly due to a lack of necessary information for long video reasoning.
Figure 7: Higher resolution (1920$\times$1080, 1280$\times$720) QA Frames generally improve performance on EgoTextVQA-Outdoor.
Q3: Are high-resolution frames for scene text critical? Yes. Most existing MLLMs are limited to accepting fixed, low-resolution images as input. When higher resolutions are used, the number of input frames is limited. We compare the use of “Low-res. video” inputs and the specified three frames for question generation as “Low-res. frames” and “High-res. frames”. Here we choose GPT-4o [35] and three open-source models that can accept frames with resolutions of up to 1920$\times$1080 and 1280$\times$720. Figure 7 shows that answering with “QA Frames” at low resolution is comparable to uniform video sampling. Yet, with the increase of resolution, GPT-4o [35], Qwen2-VL [47], InternVL2-8B [8], and MiniCPM-V 2.6 [53] all show profound improvements, up to 18.6$\sim$31.0%.
Table 6: Effects of scene-text resolution and area ratio. Scale parameter is to control the magnification ratio of the scene text regions. SR: Super Resolution Scene Text. We experiment with 30% of the data for efficiency.
  Method Scale EgoTextVQA-Outdoor EgoTextVQA-Indoor
Accuracy Score Accuracy Score
Qwen2-VL [47] - 28.2 2.0 23.3 1.8
Qwen2-VL w/ SR 1.00 $\times$ 31.9 2.1 24.2 1.8
1.25 $\times$ 33.1 2.2 23.5 1.8
1.50 $\times$ 34.1 2.2 22.3 1.7
Gemini 1.5 Pro [39] - 33.4 2.0 34.4 2.1
Gemini 1.5 Pro w/ SR 1.00 $\times$ 36.9 2.2 31.7 2.1
1.25 $\times$ 38.1 2.2 30.1 2.0
1.50 $\times$ 38.1 2.2 30.2 2.0
Table 7: Effects of different modality inputs. V: Video. Q: Question. ST: Scene Text.
  Method Input EgoTextVQA-Outdoor EgoTextVQA-Indoor
V Q ST Accuracy Score Accuracy Score
  Qwen2-VL [47] - - 2.1 0.5 2.5 0.5
- 15.4 1.2 13.3 1.1
- 28.2 2.0 23.3 1.8
39.9 $\uparrow$ 11.7 2.5 $\uparrow$ 0.5 23.6 $\uparrow$ 0.3 1.8 $\uparrow$ 0.0
LLaVA-Next-Video [58] - - 8.5 1.1 6.0 1.0
- 20.6 1.4 15.6 1.2
- 19.5 1.4 22.7 1.8
37.4 $\uparrow$ 17.9 2.3 $\uparrow$ 0.9 30.8 $\uparrow$ 8.1 2.1 $\uparrow$ 0.3
GPT-4o [35] - - 5.26 0.84 5.7 0.9
- 24.1 1.71 17.6 1.4
- 30.3 1.7 28.3 1.8
52.9 $\uparrow$ 22.6 3.0 $\uparrow$ 1.3 37.9 $\uparrow$ 9.6 2.3 $\uparrow$ 0.5
Gemini 1.5 Pro [39] - - 4.3 0.8 4.6 0.8
- 18.2 1.3 15.2 1.0
- 33.4 2.0 34.4 2.1
49.5 $\uparrow$ 16.1 2.9 $\uparrow$ 0.9 39.5 $\uparrow$ 5.1 2.4 $\uparrow$ 0.3
Aside from increasing the global frame resolution, we particularly study increasing the local scene-text resolution while keeping the frame resolution unchanged. To this end, we apply Microsoft Azure OCRfootnotetext: https://azure.microsoft.com/en-us/services/cognitive-services to detect scene text regions and then exploit the SOTA scene text super-resolution model DiffTSR [59] to increase the resolution of the detected scene text regions to 380 $\times$ 128. To project the super-resolution scene text back to the original low-resolution video frames, we consider three options: resize to 1.00 $\times$, 1.25$\times$, and 1.50$\times$ of the original scene text size. The results in Table 6 show that the higher resolution scene text effectively boosts the model performances on EgoTextVQA-Outdoor. Nevertheless, its improvements on EgoTextVQA-Indoor are unstable and the performances even decline with super-resolution scene text input. We speculate that scene-text detection in Ego4D videos is harder than in RoadTextVQA videos, due to lower resolution and challenging perspective. The lower performance improvements on EgoTextVQA-Indoor (vs. Outdoor) with additional OCR results in Table 7 further consolidate our speculation.
Q4: Is text alone enough to answer questions? No. We feed the best-performing open-source and closed-source models with only the scene text, but as Table 7 shows, none of these models can answer the questions reasonably. Pure scene text inputs are sufficient for answering some questions, but video input is more important for almost all the models. Furthermore, adding scene text together with video input further boosts the performances significantly, leading to the highest results in all our explorations. For instance, GPT-4o achieves 52.9% on EgoTextVQA-Outdoor and Gemini’s accuracy on EgoTextVQA-Indoor reaches to 39.5%. The findings demonstrate the importance of both video and scene text inputs for better QA assistance. More heuristic solution investigations can be found in Appendix B.3.
6 Conclusion
To provide a testing ground to evaluate model capacity and develop new techniques for egocentric scene-text aware QA assistance, we carefully constructed the EgoTextVQA benchmark. EgoTextVQA highlights QA assistance involving scene text from an ego-perspective in diverse real-life scenarios, including mainly outdoor driving and indoor house-keeping, with the questions reflecting the real user needs yet without the visual focus on scene text. With well-classified visual scenarios and question types, we comprehensively analyzed models that excel on existing scene-text VQA benchmarks and found that they perform poorly on EgoTextVQA. We further explore heuristic solutions and provide insights for improvements, With these efforts, we hope this work can complement existing VideoQA research in advancing real-life egocentric QA assistance.
Supplementary Material
A EgoTextVQA Dataset
A.1 Manual Participation
We present examples of QA pairs generated by GPT-4o in Figure 8 to highlight the issues of automatic generation and underscore the value of manual correction. The primary problems observed are as follows: (a) Hallucinated Answers: The generated answers are unseen from the visual environment and cannot be confirmed by the annotators. (b) Scene Text Irrelevance: The questions have vague references and fail to incorporate scene text understanding for answers. (c) Scene Text Errors: The questions or answers contain incorrect scene text. (d) and (f) Non-colloquial Questions: The questions are mechanical; they are phrased unnaturally and do not align well with daily spoken language. (e) Not Reflect User Needs: The question does not reflect real user needs and hardly occurs in human daily life. After manual participation, about 70% of the generated QAs are deleted and 30% of the remaining QAs are revised.
Figure 8: Manual participation on EgoTextVQA creation.
B Experiment
B.1 Model Details
We provide a concise introduction to the MLLMs evaluated in Section 4, as outlined below:
  • GPT-4o [35] advances the GPT-4 family towards more natural human-computer interactions.
  • Gemini 1.5 Pro [39] builds on Gemini 1.0’s [44] research advances and multimodal capabilities and it is optimized for a wide-range of reasoning tasks.
  • Gemini 1.5 Flash [39] is a model from Gemini 1.5 family offering low latency and enhanced performance.
  • Qwen2-VL [47] employs a ViT-675M [38] as the visual encoder, Qwen2-7B as the language model, and an MLP projector. It improves upon Qwen-VL [1] with (1) naive dynamic resolution, allowing ViT to handle images of varying resolutions, and (2) multimodal rotary position embedding, which decomposes positional encoding into temporal, height, and width components. Qwen2-VL is pre-trained on diverse datasets, including image-text pairs, OCR data, interleaved articles, VQA datasets, video dialogues, and image knowledge sources, enabling a stronger multimodal understanding.
  • LLaVA-NeXT-Video [58] choose SigLIP-SO400M [55] as the visual encoder, Qwen2 [47] as the language model, and a two-layer MLP as the projector. It utilizes the AnyRes [27] technique to segment high-resolution images for the visual encoder and extends this approach to video processing. LLaVA-NeXT-Video has excellent reasoning, OCR, and world knowledge capabilities, achieving strong performance in video-based multimodal tasks.
  • VILA1.5 [25] integrates CLIP-L [38] as the visual encoder, LLaMA-2 [46] as the language model, and a linear projector. It fine-tunes on a mix of internal data, including OCR-VQA [34] and ST-VQA [3], and improves contextual learning by unfreezing the LLM during interleaved image-text pre-training. VILA1.5 excels in video reasoning, in-context learning, visual chain-of-thought reasoning, and world knowledge.
  • InternVL2 [8] integrates InternViT-300M [9] with InternLM2.5 [5] via a randomly initialized MLP projector. It is trained on OCR datasets generated by PaddleOCR [22], utilizing Chinese images from Wukong and English images from LaionCOCO [40]. Building on the strong visual representations and high-resolution image processing capabilities of InternVL1.5 [8], InternVL2 incorporates instruction tuning, enabling competitive performance in document and chart comprehension, infographics QA, scene text understanding, OCR, and multimodal reasoning tasks.
  • CogVLM2-Video [18] utilizes the EVA-CLIP [43] as the visual encoder, LLaMA3-8B as the language model, and a 2$\times$2 convolutional layer followed by a SwiGLU [41] as the adapter. Unlike CogVLM [48], CogVLM2 improves pre- and post-training data diversity and quality. The Synthetic OCR Dataset, a key pre-training resource, includes four OCR scenarios: (1) synthetic OCR images with text generated in Python, (2) real-world images with PaddleOCR [22], (3) academic papers with extracted LaTeX via Nougat [4], and (4) HTML/LaTeX-rendered tables and formulae. CogVLM2-Video adapts CogVLM2 for videos, enhancing open-domain QA with temporal localization and timestamp-aware QA.
  • MiniCPM-V 2.6 [53] employs SigLIP-SO400M [55] as the visual encoder, Qwen2 [47] as the language model, and a compression module with one-layer cross-attention and a moderate number of queries as the projector. Its training includes pre-training on English and Chinese image captioning and OCR data, followed by fine-tuning on datasets like TextVQA [42], OCR-VQA [34], and ST-VQA [3]. MiniCPM-V 2.6 excels in conversational and reasoning tasks across multiple images and videos, with high-resolution perception enabling features like table-to-markdown conversion and OCR transcription.
  • ShareGPT4Video [7] builts on LLaVA-Next-8B [21]. Based on the proposed ShareGPT4Video [7] dataset, the proposed captioning model ShareCaptioner-Video generates high-quality captions with detailed temporal descriptions for various videos. ShareCaptioner-Video is fine-tuned with the collected video caption data. For video understanding, ShareGPT4Video’s training dataset combines VQA samples from various instructional video-to-text datasets with video-caption pairs.
B.2 Study of MLLM Design
As the results in Tables 3 and 4, we analyze key factors contributing to the superior performance of strong models (Qwen2-VL [47], LLaVA-NeXT-Video [58], and InternVL2 [9]): (1) enhanced visual encoder capable of handling high-resolution and long-video inputs. As shown in Table 8, increasing the number of video frames and resolution improves Qwen2-VL’s performance by 1.2% and 5.8%, respectively.; (2) more powerful LLM backbones. Compared with InternVL2-8B, InternVL2-26B performance has a 7% increases; and (3) large-scale OCR training data. Beyond the commonly used TextVQA datasets, InternVL2 leverages PaddleOCR to generate additional OCR training samples. Additionally, we observe that as the number of video frames increases, Qwen2-VL’s performance improves, whereas InternVL2’s declines, underscoring the effectiveness of Qwen2-VL’s video embedding design.
B.3 Heuristic Solution Investigations
Effect of Timestamp-Aware Sampling We further investigate an alternative question-timestamp aware sampling strategy: starting from the question timestamp and uniformly sampling within fixed durations of 4 seconds and 32 seconds. As shown in Table 9, on EgoTextVQA-Outdoor, when sampling the same number of frames (#F=16), this fixed duration sampling strategy achieves comparable or even superior performance to standard uniform sampling across the whole video for both Qwen2-VL [47] and Gemini 1.5 Pro [39]. However, on EgoTextVQA-Indoor, when sampling #F=48, we observe that while Qwen2-VL [47] maintains comparable performance to standard uniform sampling, the performance of Gemini 1.5 Pro [39] drops by by about 7%. This decline may stem from Gemini 1.5 Pro’s stronger performance on questions requiring long-term video comprehension, which is less effectively captured by this fixed-duration sampling approach.
Combination of Heuristic Strategies In the main text, we have explored different heuristic strategies separately. Here, we additionally study the combinations of heuristic strategies. First, for timestamp-aware video sampling, we adopt the strategy of “fixed-duration sampling” to EgoTextVQA-Outdoor and “1fps-backward sampling” to EgoTextVQA-Indoor, inspired by the results of the two different video sampling strategies on these two datasets. The results in Table 10 show that the models achieve cumulative performance improvements as heuristic strategies are progressively applied on EgoTextVQA-Outdoor. Yet, the improvements are not stable on EgoTextVQA-Indoor, suggesting the significant challenge of egocentric scene-text aware QA assistance in daily house-keeping.
Table 8: Study of MLLM design on EgoTextVQA-Outdoor. The alignment module is MLP layer. VE: Visual Encoder.
Method VE Res. #F Accuracy Score
Qwen2-VL [47] ViT-675M $448^{2}$ 16 22.4 1.6
ViT-675M $448^{2}$ 32 23.6 1.7
ViT-675M - 16 28.2 2.0
InternVL2-8B [9] InternViT-300M $448^{2}$ 16 16.5 1.3
InternViT-300M $448^{2}$ 32 16.4 1.3
InternVL2-26B [9] InternViT-6B $448^{2}$ 16 23.5 1.7
Table 9: Effects of different numbers of video frames uniformly sampled within fix-duration before the question timestamp. We set a fixed duration of 4 seconds in EgoTextVQA-Outdoor and 32 seconds in EgoTextVQA-Indoor. S: Standard Uniformly Sampling. F: Fix-Duration Sampling. We experiment with 30% of the data for efficiency.
  Method EgoTextVQA-Outdoor EgoTextVQA-Indoor
#F Accuracy Score #F Accuracy Score
  Qwen2-VL [47] w/ S 16 28.2 2.0 48 23.3 1.8
Qwen2-VL [47] w/ F 4 26.1 1.8 12 16.2 1.4
8 29.3 2.0 24 18.3 1.5
12 29.9 2.0 36 22.1 1.7
16 30.9 2.1 48 22.7 1.7
Gemini 1.5 Pro [39] w/ S 32 33.4 2.0 60 34.4 2.1
Gemini 1.5 Pro [39] w/ F 4 27.9 1.7 12 22.8 1.6
8 32.9 1.9 24 27.1 1.7
12 33.5 2.0 36 29.0 1.8
16 33.4 2.0 48 27.3 1.8
B.4 Human Study
The human results in Tables 3 and 4 are based on two rounds of standard human studies. Based on the reason analysis for the poor human performance in Section 4.2, we further validate the human performance by reducing the scene text recognition challenge. We sample 100 additional questions for humans to answer by providing the corresponding question frames. Table 11 shows that humans perform better without the challenge of temporal grounding but still lag behind the best closed-source model (GPT-4o [35]). This suggests advanced models may surpass humans in scene-text recognition or external knowledge, highlighting the importance of research on scene-text QA assistance.
Table 10: Effects of combining different heuristic strategies. T: Timestamp-Aware Sampling. ST: Additional Scene Text Input. HR: High-Resolution Scene Text (Scale = 1.25$\times$). We experiment with 30% of the data for efficiency.
  Method Input EgoTextVQA-Outdoor EgoTextVQA-Indoor
T ST HR Accuracy Score Accuracy Score
  Qwen2-VL [47] - - - 28.2 2.0 23.3 1.8
- - 30.9 2.1 22.6 1.7
- 42.4 2.7 25.3 1.8
42.6 2.7 27.1 1.9
Gemini 1.5 Pro [39] - - - 33.4 2.0 34.4 2.1
- - 34.7 2.0 31.1 2.0
- 49.5 2.9 38.0 2.3
51.1 3.0 36.5 2.2
Table 11: Results of using video vs. QA frames (three frames for QA generation) on EgoTextVQA-Indoor.
Method Video QA Frames
Accuracy Score Accuracy Score
Human 26.0 1.9 36.0 2.3
GPT-4o [27] 25.0 1.6 39.0 2.4
Gemini-1.5 Pro [29] 33.0 2.0 35.0 2.2
Figure 9: Result visualization on EgoTextVQA.
B.5 Case Analysis
As shown in Figure 9, we qualitatively analyze the performance of MLLMs on EgoTextVQA. For EgoTextVQA-Outdoor, the “Description” example shows that all models struggle to accurately identify the target referred to by the question at the queried timestamp. The “Location” example shows that only Gemini 1.5 Pro [39] correctly inferred the intention of the question and provided the precise answer. For the real-time “Direction” examples, where the same question is posed at different timestamps, which corresponds to different answers, the left example shows that the questioned building is located right in front of the user at the question timestamp (4.3s), but all models fail to provide the correct answer, likely due to poor 3D spatial relation reasoning. Also, if the user poses the same question at 7.2s when he has moved to the left of the building, all models are unresponsive to such visual changes and tend to keep their original answers. This indicates that the models struggle to provide reasonable answers based on the real-time visual context in dynamic environment.
For EgoTextVQA-Indoor, the “Shopping” example shows that all models fail to effectively answer the total expense of 46.85 after the checkout process, indicating their limitations in infer the total number after observing the price changes on the cashier’s display during the checkout process. In the “Book” example, the models need to identify the book from a large collection of books that matches the user’s needs (e.g., related to topics on plants or nature). The wrong answers indicate that the corresponding models are either weak at scene-text recognition and knowledge reasoning or tend to hallucinate their responses to match some question key words. Similar issues are also observed in the “Hands-on” example. Finally, the failures in the “Gameplay” example suggest that most models are weak in reasoning the real-time object state and people’s real-time actions from an ego point of view. For example, while the “rule sheet” is on the table most of the time, it is on the other game player’s hand at the time of user questioning.
C Agreement between Human and Evaluator
In this section, we evaluate the performance of models on EgoTextVQA-Outdoor using GPT-4o mini [36] and human annotators. Following [11], we invite three annotators to assess GPT-4o [35] and Gemini 1.5 Pro [39], the overall best-performing model. Human annotators maintain the same scoring principle as the model, as shown in Table 16. We randomly sample 100 QA pairs for evaluation. As shown in Table 12, GPT-4o mini and human annotators achieve similar Accuracy and Score, with Pearson correlation coefficients of 0.80 and 0.87, respectively, indicating strong consistency. The Cohen’s Kappa coefficients among three volunteers are 0.77 on Accuracy, indicating a high human agreement. To ensure reproducibility, future evaluations should use the same model version (GPT-4o-mini-2024-07-18) and the prompt in Table 16.
Table 12: Judgments of human and GPT-4o mini.
Method GPT-4o [35] Gemini 1.5 Pro [39]
Accuracy Score Accuracy Score
Human 36.0 1.9 47.3 2.5
GPT-4o mini [36] 34.0 1.8 42.0 2.3
D Model Prompts
Table 13 provides the prompts used by GPT-4o for question generation and filtering. Table 14 lists the prompts employed by GPT-4o for automatic question label annotation. Table 15 details the specific prompts applied for model inference. Table 16 shows the prompts used by GPT-4o-mini for model evaluation. Table 17 includes the prompts designed for heuristic solutions with different modality inputs.
Table 13: Prompts for question-answer generation and filtering on EgoTextVQA
Question-Answer Generation Prompts
## Question Prompt:
Give you a first-person perspective video, which records the scene you see from the first-person perspective. Please judge from your perspective whether there are scene texts in the video. If so, please tell me what these scene texts are. Then, you have some questions about the scene texts you see, and ask three questions related to the activities you are going to carry out. These scene texts can serve as clues to help you answer your questions. Please generate three highly diverse questions based on the scene texts related to your activities in the first-person perspective video. If there is no scene text in the video, it is not necessary. Your questions should meet the following requirements:
Requirement 1: The questions should involve scene text understanding in the video.
Requirement 2: The questions should be goal-oriented and relevant to human daily life.
Requirement 3: The questions should require understanding multiple video frames, not just a single frame.
Requirement 4: The questions should be asked from a first-person perspective, expressed as colloquially as possible, and the first-person pronoun “I” should be used appropriately.
Requirement 5: The questions should be of moderate length.
When announcing the question please label each question as “Question 1, 2, 3: {question}”.
Please start your questions with the question word “what”, “where”, “which”, etc. You don’t need to explain too much about what you are doing or indicate the location of the scene text in the video. Avoid the words “video” and “frame” in the questions. Remember to make sure that the correct answer to your question can be taken directly from the video and is concise enough.
Examples of good questions:
“Question 1: Which way is the exit?”
“Question 2: Could you tell me how much this item costs?”
“Question 3: What is the speed limit on this road?”
Image:{image1} Image:{image2} Image:{image3}
## Answer Prompt:
I provide three questions as follows: {question}
You need to create an exam that tests above student abilities based on the three questions I just provided. Each question should have open-ended but short correct answers. Your answers have the following requirements:
Requirement 1: Your answers should be short and be closely related to the scene text in the video.
Requirement 2: Your answers should not mention any particular video frame number.
Requirement 3: Do not use letters for the answer choices.
You must print one correct answer and four wrong answers on separate lines in the following format: ¨
Correct Answer :{answer}
Automatic Filtering Prompt
You are a helpful assistant. You can answer the following questions based on your general knowledge.
Question: {question} Answer briefly with a single word, a phrase, or a short sentence.
Table 14: Prompts for GPT-4o to annotate question categories on EgoTextVQA-Outdoor.
Question Classification Prompts
Question: {question}
Which of the following five categories does this question belong to? Please only answer the category name, such as Direction.
1. Location: Questions about a place or location. For example:
1) Where is the gas station?
2) Which stores can I find on the right side of the road at this intersection?
2. Direction: Questions related to navigation, driving direction, and turns. For example:
1) Is the next road a left or right turn?
2) If I want to go to Cava, on which side of the street should I look for it?
3) Where should trucks go according to the signs?
3. Description: Questions that focus on scene text such as road signs, price labels, and billboards. For example:
1) What does the sign on the side of the road say?
2) What is the name of the center on the left side of the road?
3) What is the name of the street to my right?
4. Intention Reasoning: Questions about behavioral activities involving drivers or passengers to solve personal needs. For example:
1) Where do I need to go to solve my financial problems?
2) Is there a place nearby where I can shop for appliances and electronics?
5. Others: Composite questions that involve multiple different or the same types of the above, such as asking about both description and location. For example:
1) What event is being advertised on the bus, and where is it taking place?
2) What is the contact number for the leadspace building, and what service might they provide?
Table 15: Prompts for MLLM inference on EgoTextVQA.
Model General Prompts
GPT-4o Based on the following images from a video, please briefly answer the following question with a single word, a phrase, or a short sentence. Question: {question}. Output the answer to the question in the following format: Answer: {answer}. If you cannot answer the question, please answer “Unanswerable” and briefly explain why you cannot answer.
Gemini 1.5 Flash Based on the following images from a video, please briefly answer the following question with a single word, a phrase, or a short sentence. Question: {question}. Output the answer to the question in the following format: Answer: {answer}. If you cannot answer the question, please answer “Unanswerable” and briefly explain why you cannot answer.
Gemini 1.5 Pro Based on the following images from a video, please briefly answer the following question with a single word, a phrase, or a short sentence. Question: {question}. Output the answer to the question in the following format: Answer: {answer}. If you cannot answer the question, please answer “Unanswerable” and briefly explain why you cannot answer.
LLaVA-Next-Video Please answer the following questions related to this video. If you cannot answer the question, please answer “Unanswerable” and briefly explain why you cannot answer. Keep your answer as short as possible. Keep your answer as short as possible. Keep your answer as short as possible. Question: {question}
CogVLM2-Video You are a person in the situation shown in the following consecutive images from a video. You can answer questions that humans ask to help them make decisions. Now you are observing your surroundings and answering questions based on the current situation. Understanding the scene text around you is important for answering questions. Answer the questions in the first-person perspective. If you cannot answer the question, please answer “Unanswerable” and briefly explain why you cannot answer. Question: {question}
InternVL2 You are a person in the situation shown in the following consecutive images from a video. You can answer questions that humans ask to help them make decisions. Now you are observing your surroundings and answering questions based on the current situation. Understanding the scene text around you is important for answering questions. Answer the questions in the first-person perspective. If you cannot answer the question, please answer ’Unanswerable’ and briefly explain why you cannot answer. Keep your answer as short as possible! Keep your answer as short as possible! Keep your answer as short as possible! Question: {question}
Qwen2-VL You are a person in the situation shown in the following consecutive images from a video. You can answer questions that humans ask to help them make decisions. Now you are observing your surroundings and answering questions based on the current situation. Understanding the scene text around you is important for answering questions. Answer the questions in the first-person perspective. If you cannot answer the question, please answer ’Unanswerable’ and briefly explain why you cannot answer. Question: {question}
VILA1.5 You are a helpful language and vision assistant. You are able to understand the visual content that the user provides, and assist the user with a variety of tasks using natural language. Question: {question}
ShareGPT4Video You are a person in the situation shown in the following consecutive images from a video. You can answer questions that humans ask to help them make decisions. Now you are observing your surroundings and answering questions based on the current situation. Understanding the scene text around you is important for answering questions. Answer the questions in the first-person perspective. If you cannot answer the question. please answer “Unanswerable” and briefly explain why you cannot answer. Question: {question}
MiniCPM-V 2.6 You are a person in the situation shown in the following consecutive images from a video. You can answer questions that humans ask to help them make decisions. Now you are observing your surroundings and answering questions based on the current situation. Understanding the scene text around you is important for answering questions. Answer the questions in the first-person perspective. If you cannot answer the question, please answer ’Unanswerable’ and briefly explain why you cannot answer. Keep your answer as short as possible! Keep your answer as short as possible! Keep your answer as short as possible! Question: {question}
Table 16: Prompts for GPT-4o-mini to evaluate MLLMs on EgoTextVQA.
Evaluation Prompts
You are an intelligent chatbot designed for evaluating the correctness of generative outputs for question-answer pairs. Your task is to compare the predicted answer with the correct answer and determine if they match meaningfully. Here’s how you can accomplish the task:
##INSTRUCTIONS:
- Focus on the meaningful match between the predicted answer and the correct answer. Please note that not only matches of noun phrases between answers, but also matches of prepositional phrases.
For example, “at the car wash on your right” does not exactly match “car wash”. “at the gas station beside the sign ’gas sale’” does not exactly match “gas station””
- Consider synonyms or paraphrases as valid matches. Note that the predicted answer must be consistent with the string type of the correct answer, which may include phone numbers, email addresses, numbers, dates, etc.
For example, the string types “www.usps.com” and “visit their website” are inconsistent, the string types “9849041316” and “advertiser’s contact number” are inconsistent.”
- Evaluate the correctness of the prediction compared to the answer.”
Please evaluate the following video-based question-answer pair:
Question: {question} Correct Answer: {GT answer} Predicted Answer: {predicted answer}
Provide your $eval_{code}$ only as a yes/no and score where the score is an integer value between 0 and 5, with 5 indicating the highest meaningful match.
Please generate the response in the form of a Python dictionary string with keys ’pred’ and ’score’, where the value of ’pred’ is a string of ’yes’ or ’no’ and the value of ’score’ is in INTEGER, not STRING. DO NOT PROVIDE ANY OTHER OUTPUT TEXT OR EXPLANATION. Only provide the Python dictionary string. For example, your response should look like this: {’pred’: ’yes’, ’score’: 5}, {’pred’: ’no’, ’score’: 1}.
Table 17: Prompts for heuristic solution study of different modality inputs on EgoTextVQA.
Model Input Prompts
w/ Q You are a helpful assistant. You can answer the following questions based on your general knowledge. Question: {question}
w/ Q & ST You are a helpful assistant. You are provided with some important scene text information. You can answer the following questions based on your common sense or the scene text information I provide. Please answer as briefly as possible. Please note that this scenario text information is very important. You can find the scene text related to the question as the answer. Scene Text: {OCR results} Question: {question}
w/ V & Q & ST You are a person in the situation shown in the following consecutive images from a video. You can answer questions that humans ask to help them make decisions. Now you are observing your surroundings and answering questions based on the current situation. I will provide you with the following scene text that may be included in each image. Understanding the scene text is important for answering questions. Answer the questions in the first-person perspective. If you cannot answer the question, please answer ’Unanswerable’ and briefly explain why you cannot answer. The scene texts in Frame 0 include: {OCR results}. The scene texts in Frame 1 include: {OCR results}. The scene texts in Frame 2 include: {OCR results}. The scene texts in Frame {frame id} include: {OCR results}. Question: {question}